OpenStudy (anonymous):

Medal Solve for x. Give exact answers in ascending order supposing a is positive. 4x2 + 24ax + 20a2 = 0

OpenStudy (31356):

Is there any answer choices?

OpenStudy (anonymous):

no sorry

Directrix (directrix):

4x2 + 24ax + 20a2 = 0 Divide each term by 4. x^2 + 6ax + 5 a^2 = 0 Factor (x + 5a) (x + 1a) = 0 Zero Product Property x + 5a = 0 or x + 1a = 0 x = - 5a or x = -a ascending order -- small to larger -5a, -a @harz360